# Data-driven Implementations of Various Generalizations of Balanced Truncation

### v1

@@ Abstract
There exist two main frameworks for non-intrusive implementations of approximate balanced truncation: the quadrature-based framework [1] and the ADI-based framework [2]. Both approaches rely solely on samples of the transfer function to construct truncated balanced models, eliminating the need for access to the original model’s state-space realization. Recently, the quadrature-based framework has been extended to various generalizations of balanced truncation, including positive-real balanced truncation, bounded-real balanced truncation, and balanced stochastic truncation. While this extension [3] is theoretically non-intrusive—meaning it does not require the original state-space realization—it depends on samples of spectral factorizations of the transfer function. Since practical methods for obtaining such samples are currently unavailable, this extension remains largely a theoretical contribution.
In this work, we present a non-intrusive ADI-type framework for these generalized balanced truncation methods that requires only samples of the original transfer function for implementation. To achieve this, we first demonstrate how the approximations of ADI-type method for Riccati equations (RADI) [4] can be equivalently obtained via projection-based interpolation. Leveraging this insight, we develop several non-intrusive, data-driven algorithms that not only satisfy the desired interpolation conditions but also preserve key system properties such as positive-realness, bounded-realness, and minimum phase. Notably, these algorithms rely solely on solving Lyapunov and Sylvester equations, avoiding computationally intensive Riccati equations or linear matrix inequalities. Additionally, we propose non-intrusive, data-driven implementations of RADI (and ADI)-based generalizations of balanced truncation, covering LQG balanced truncation, $\mathcal{H}_{\infty}$ balanced truncation, positive-real balanced truncation, bounded-real balanced truncation, self-weighted balanced truncation, and balanced stochastic truncation. We also discuss practical considerations for improving computational efficiency. Numerical experiments demonstrate that the proposed algorithms perform comparably to their intrusive counterparts.
@@ 1 Preliminaries
with matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, $C\in\mathbb{R}^{p\times m}$, and $D\in\mathbb{R}^{p\times m}$.
@@
The reduced-order model ROM $\hat{G}(s)$ is constructed from $G(s)$ through Petrov-Galerkin projection:
@@
using full-rank projection matrices $\hat{V},\hat{W}\in\mathbb{C}^{n\times r}$ satisfying $\hat{W}^{*}\hat{V}=I$. Importantly, the ROM remains unchanged under transformations $\hat{V}\rightarrow\hat{V}T_{v}$ and $\hat{W}\rightarrow\hat{W}T_{w}$, where $T_{v},T_{w}\in\mathbb{C}^{r\times r}$ are nonsingular. This property allows the complex projection matrices $(\hat{V},\hat{W})$ and corresponding reduced-order state-space matrices $(\hat{A},\hat{B},\hat{C})$ to be converted to real-valued representations through appropriate similarity transformations [5].
@@ 1.2 Interpolatory Loewner framework [7]
$\hat{W}^{*}B$ $=\begin{bmatrix}H(\mu_{1})\\ \vdots\\ H(\mu_{w})\end{bmatrix},\quad\hat{C}=C\hat{V}=\begin{bmatrix}H(\sigma_{1})&\cdots H(\sigma_{v})\end{bmatrix},$ (5)
@@
1.3 Pseudo-optimal Rational Krylov Algorithm [8]
@@
$S_{v}$ $=\text{diag}(\sigma_{1},\dots,\sigma_{r})\otimes I_{m},$ $S_{w}$ $=\text{diag}(\mu_{1},\dots,\mu_{r})\otimes I_{m},$
$L_{v}$ $=\begin{bmatrix}1,\dots,1\end{bmatrix}\otimes I_{m},$ $L_{w}^{T}$ $=\begin{bmatrix}1,\dots,1\end{bmatrix}\otimes I_{m}.$ (6)
@@
meets the $\mathcal{H}_{2}$-optimality condition $\frac{\partial}{\partial\hat{C}}\big{(}||G(s)-\hat{G}(s)||_{\mathcal{H}_{2}}^{2}\big{)}$. This method is referred to as Input PORK (I-PORK) throughout this paper.
@@
satisfies the $\mathcal{H}_{2}$-optimality condition given by $\frac{\partial}{\partial\hat{B}}\big{(}\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{2}}^{2}\big{)}$. This approach is referred to as Output PORK (O-PORK) throughout this paper.
1.4 Balanced Truncation Family
@@
[…]
@@
The balanced square-root algorithm (BSA) [9] is among the most numerically stable methods for implementing balanced truncation (BT) [10]. It begins by computing the singular value decomposition (SVD) of the product $L_{q}^{T}L_{p}$:
@@
In Linear Quadratic Gaussian (LQG) BT (LQG-BT) [11], the Gramian-like matrices are computed by solving the following filter and controller Riccati equations:
@@
$A^{T}Q_{\text{LQG}}+Q_{\text{LQG}}A^{T}+C^{T}C-Q_{\text{LQG}}BB^{T}Q_{\text{LQG}}=0.$ (17)
@@
In $\mathcal{H}_{\infty}$ BT ($\mathcal{H}_{\infty}$-BT) [12], the Gramian-like matrices are computed by solving the following filter and controller Riccati equations with $\gamma>0$:
@@
$A^{T}Q_{\mathcal{H}_{\infty}}+Q_{\mathcal{H}_{\infty}}A^{T}+C^{T}C-(1-\gamma^{-2})Q_{\mathcal{H}_{\infty}}BB^{T}Q_{\mathcal{H}_{\infty}}=0.$ (19)
@@
In positive-real BT (PR-BT) [13, 14], the Gramians solve the following Riccati equations:
@@
In bounded-real BT (BR-BT) [15, 14], the Gramians solve the following Riccati equations:
@@
In self-weighted BT (SW-BT) [16], the controllability Gramian $P$ remains the same as in BT, while the weighted observability Gramian $Q_{\text{SW}}$ solves the following Lyapunov equation:
@@
In balanced stochastic truncation (BST) [13, 17], the controllability Gramian $P$ remains the same as in BT, while the weighted observability Gramian $Q_{\text{S}}$ solves the following Riccati equation:
@@
1.5 Quadrature-based Non-intrusive Implementation of Balanced Truncation
@@
$P$ $\approx\tilde{P}=\sum_{i=1}^{n_{p}}w_{p,i}^{2}(j\omega_{i}I-A)^{-1}BB^{T}(-j\omega_{i}I-A^{T})^{-1},$
$Q$ $\approx\tilde{Q}=\sum_{i=1}^{n_{q}}w_{q,i}^{2}(-j\nu_{i}I-A^{T})^{-1}C^{T}C(j\nu_{i}I-A)^{-1},$
where $\omega_{i}$ and $\nu_{i}$ are the quadrature nodes, and $w_{p,i}^{2}$ and $w_{q,i}^{2}$ are their corresponding weights. The factorizations $\tilde{P}=\tilde{L}_{p}\tilde{L}_{p}^{T}$ and $\tilde{Q}=\tilde{L}_{q}\tilde{L}_{q}^{T}$ can be expressed as: $\tilde{L}_{p}=\hat{V}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}\hat{L}_{q}$, with
$\hat{V}$ $=\begin{bmatrix}(j\omega_{1}I-A)^{-1}B&\cdots&(j\omega_{n_{p}}I-A)^{-1}B\end{bmatrix},$ (26)
$\hat{W}$ $=\begin{bmatrix}(-j\nu_{1}I-A^{T})^{-1}C^{T}&\cdots&(-j\nu_{n_{q}}I-A^{T})^{-1}C^{T}\end{bmatrix},$ (27)
$\hat{L}_{p}$ $=\text{diag}(w_{p,1},\dots,w_{p,n_{p}})\otimes I_{m},$
$\hat{L}_{q}$ $=\text{diag}(w_{q,1},\dots,w_{q,n_{q}})\otimes I_{m}.$
@@
Finally, the ROM in Quadrature-based BT (QuadBT) [1] is constructed as:
@@
    Practical methods to obtain these samples do not currently exist and remain as future work.
@@
1.6 ADI-based Non-intrusive Implementation of Balanced Truncation
@@ 2 Main Work
In this section, we present several new results on data-driven model reduction and controller design, outlined as follows:
@@ 2.1 RADI-based Non-intrusive Implementation of LQG-BT
    Given $\hat{A}=\hat{P}_{\text{LQG}}(-S_{v}^{*}+\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C})\hat{P}_{\text{LQG}}^{-1}$, the eigenvalues of $\hat{A}$ are equal to those of $-S_{v}^{*}+\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$. As $-S_{v}^{*}$ is Hurwitz, the matrix $\hat{A}-\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$ is Hurwitz with eigenvalues at $(-\sigma_{1}^{*},\cdots,-\sigma_{v}^{*})$.
@@
Let $\hat{W}$ be as defined in (2), with all interpolation points $(\mu_{1},\dots,\mu_{w})$ located in the right half of the $s$-plane. Assume further that the pair $(S_{w},L_{w})$ is controllable and that $P_{v}>0$ uniquely solves the Lyapunov equation:
@@
Similar to the ADI-based non-intrusive implementation of BT, the RADI-based non-intrusive implementation of LQG-BT can be directly obtained using the results of Theorems 2.1 and 2.2. The pseudo-code for this implementation, named “Data-driven LQG-BT (DD-LQG-BT)”, is provided in Algorithm 1.
Algorithm 1 DD-LQG-BT
@@
Output: ROM: $(\hat{A},\hat{B},\hat{C})$
1:  Compute the Loewner quadruplet $(\hat{W}^{*}\hat{V},\hat{W}^{*}A\hat{V},\hat{W}^{*}B,C\hat{V})$ from the samples $H(\sigma_{i})$ and $H(\mu_{i})$ via (5).

### v2

@@ Abstract
There exist two main approaches for non-intrusive implementations of approximate balanced truncation within the Loewner framework: the quadrature-based method [1] and the Alternating Direction Implicit (ADI)-based method [2]. Both approaches rely solely on samples of the transfer function to construct truncated balanced models, eliminating the need for access to the original model’s state-space realization. Recently, the quadrature-based approach has been extended to various generalizations of balanced truncation, including positive-real balanced truncation, bounded-real balanced truncation, and balanced stochastic truncation. While this extension [3] is theoretically non-intrusive—meaning it does not require the original state-space realization—it depends on samples of spectral factorizations of the transfer function. Since practical methods for obtaining such samples are currently unavailable, this extension remains largely a theoretical contribution.
In this work, we present a non-intrusive ADI-type framework for these generalized balanced truncation methods that requires only samples of the original transfer function for implementation. To achieve this, we first demonstrate how the approximations of ADI-type method for Riccati equations (RADI) [4] can be equivalently obtained via projection-based interpolation. Leveraging this insight, we develop several non-intrusive, data-driven algorithms that not only satisfy the desired interpolation conditions but also preserve key system properties such as positive-realness, bounded-realness, and minimum phase. Notably, these algorithms rely solely on solving Lyapunov and Sylvester equations, avoiding computationally intensive Riccati equations or linear matrix inequalities. Additionally, we propose non-intrusive, data-driven implementations of RADI (and ADI)-based generalizations of balanced truncation, covering Linear Quadratic Gaussian (LQG) balanced truncation, $\mathcal{H}_{\infty}$ balanced truncation, positive-real balanced truncation, bounded-real balanced truncation, self-weighted balanced truncation, and balanced stochastic truncation. We also discuss practical considerations for improving computational efficiency. Numerical experiments demonstrate that the proposed algorithms perform comparably to their intrusive counterparts.
@@
1 Introduction
Model order reduction (MOR) refers to system-theoretic techniques used to obtain simplified models that retain the essential input-output behavior of high-dimensional dynamical systems. These reduced order models (ROMs) are constructed to preserve the key dynamics and properties of the original system while being much lower in dimension. As a result, they offer significant computational advantages, making simulation, control, and analysis more efficient. For a detailed overview of MOR methods, refer to [5, 6, 7, 8].
Among MOR techniques, balanced truncation (BT) is a widely adopted and effective approach for linear systems [9]. It ensures that the ROM preserves the stability of the original system and provides guaranteed a priori error bound. BT truncates states that contribute minimally to the system’s input-output behavior, as determined by the system’s Hankel singular values. This results in a ROM that closely approximates the full-order model.
The main computational burden in BT lies in solving high-order Lyapunov equations to obtain the system Gramians. Numerous efficient algorithms, as discussed in surveys [10, 11], have been proposed for this purpose. These approaches require access to the system’s full state-space model and are therefore considered intrusive. In contrast, non-intrusive methods do not rely on internal state-space information but instead use system response data, such as transfer function samples or impulse responses [12, 13, 14, 15]. A notable example is the Quadrature-BT (QuadBT) algorithm introduced in [1], which constructs the ROM using numerical integration based on transfer function evaluations along the $j\omega$-axis or on sampled impulse responses and their derivatives. Over the last twenty years, the low-rank Alternating Direction Implicit (ADI) method has become a powerful tool for reducing the computational cost of BT in large-scale settings [16]. Recently, a data-driven, non-intrusive implementation of the ADI-based BT, which constructs the ROM using transfer function samples taken from the right half of the s-plane, was proposed in [2].
BT has been generalized to various classes of dynamical systems [17], including descriptor systems [18], parametric systems [19], nonlinear systems [20], and second-order systems [21], while maintaining important properties such as passivity [22], contractivity [23], and the minimum-phase condition [24]. Recently, QuadBT was extended to include certain generalizations of BT, such as balanced stochastic truncation (BST) [24], positive-real BT (PR-BT) [24], and bounded-real BT (BR-BT) [23, 25]. Although this extension [3] is theoretically non-intrusive, it relies on access to samples of spectral factorizations of the transfer function. As practical techniques for acquiring such samples are not yet available, the method remains primarily of theoretical interest.
In this work, we introduce non-intrusive, data-driven interpolation algorithms that not only interpolate the original transfer function at chosen interpolation points but also preserve key system properties such as positive-realness, bounded-realness, and the minimum-phase condition. Additionally, we present an ADI-based non-intrusive data-driven framework applicable to several BT generalizations, including Linear Quadratic Gaussian-BT (LQG-BT) [26], $\mathcal{H}_{\infty}$-BT [27], PR-BT [24], BR-BT [23], self-weighted BT (SW-BT) [28], and BST [24]. In contrast to the quadrature-based approach in [3], the proposed methods rely on transfer function samples that can be practically obtained using frequency-domain techniques from the system identification literature [29]. The data-driven algorithms for LQG-BT and $\mathcal{H}_{\infty}$-BT proposed in this paper provide a control design approach that can be used to design LQG and $\mathcal{H}_{\infty}$-infinity controllers for unknown plants using only samples of the plant’s transfer function. Numerical results using a benchmark test model demonstrate that the proposed algorithms perform comparably to their intrusive counterparts.
The paper is structured as follows. Section 2 introduces the MOR problem and provides a brief overview of key MOR techniques relevant to this work. Section 3 presents the main contributions. In Subsection 3.1, we propose non-intrusive data-driven implementations of LQG-BT and $\mathcal{H}_{\infty}$-BT. Subsection 3.2 introduces a data-driven implementation of PR-BT. Subsection 3.3 focuses on a data-driven BR-BT implementation. In Subsection 3.4, we propose a data-driven implementation of SW-BT. Subsection 3.5 presents a data-driven implementation of BST. Section 4 addresses practical considerations to prevent numerical issues during implementation. Section 5 evaluates the performance of the proposed algorithms through a benchmark numerical example. Finally, Section 6 offers concluding remarks.
@@ 2 Preliminaries
with matrices $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times m}$, $C\in\mathbb{R}^{p\times n}$, and $D\in\mathbb{R}^{p\times m}$.
@@
The ROM $\hat{G}(s)$ is constructed from $G(s)$ through Petrov-Galerkin projection:
@@
[…]
@@ 2.2 Interpolatory Loewner framework [12]
$\hat{W}^{*}B$ $=\begin{bmatrix}H(\mu_{1})\\ \vdots\\ H(\mu_{w})\end{bmatrix},\quad C\hat{V}=\begin{bmatrix}H(\sigma_{1})&\cdots H(\sigma_{v})\end{bmatrix}$, (5)
@@
2.3 Pseudo-optimal Rational Krylov Algorithm (PORK) [32]
@@
$S_{v}$ $=\text{diag}(\sigma_{1},\dots,\sigma_{v})\otimes I_{m},$ $S_{w}$ $=\text{diag}(\mu_{1},\dots,\mu_{w})\otimes I_{p},$
$L_{v}$ $=\begin{bmatrix}1,\dots,1\end{bmatrix}\otimes I_{m},$ $L_{w}^{T}$ $=\begin{bmatrix}1,\dots,1\end{bmatrix}\otimes I_{p}.$ (6)
@@
[…]
@@
satisfies the $\mathcal{H}_{2}$-optimality condition given by $\frac{\partial}{\partial\hat{B}}\big(\|G(s)-\hat{G}(s)\|_{\mathcal{H}_{2}}^{2}\big)=0$. This approach is referred to as Output PORK (O-PORK) throughout this paper.
2.4 Balanced Truncation (BT) Family
@@
$P$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}(j\omega I-A)^{-1}BB^{T}(-j\omega I-A^{T})^{-1}\,d\omega,$ (12)
$Q$ $=\frac{1}{2\pi}\int_{-\infty}^{\infty}(-j\omega I-A^{T})^{-1}C^{T}C(j\omega I-A)^{-1}\,d\omega.$ (13)
@@
The balanced square-root algorithm (BSA) [33] is among the most numerically stable methods for implementing BT [9]. It begins by computing the singular value decomposition (SVD) of the product $L_{q}^{T}L_{p}$:
@@
In LQG-BT [26], the Gramian-like matrices are computed by solving the following filter and controller Riccati equations:
@@
$A^{T}Q_{\text{LQG}}+Q_{\text{LQG}}A+C^{T}C-Q_{\text{LQG}}BB^{T}Q_{\text{LQG}}=0.$ (17)
@@
In $\mathcal{H}_{\infty}$-BT [27], the Gramian-like matrices are computed by solving the following filter and controller Riccati equations with $\gamma>0$:
@@
$A^{T}Q_{\mathcal{H}_{\infty}}+Q_{\mathcal{H}_{\infty}}A+C^{T}C-(1-\gamma^{-2})Q_{\mathcal{H}_{\infty}}BB^{T}Q_{\mathcal{H}_{\infty}}=0.$ (19)
@@
In PR-BT [24, 34], the Gramians solve the following Riccati equations:
@@
In BR-BT [35, 34], the Gramians solve the following Riccati equations:
@@
In SW-BT [28], the controllability Gramian $P$ remains the same as in BT, while the weighted observability Gramian $Q_{\text{SW}}$ solves the following Lyapunov equation:
@@
In BST [24, 36], the controllability Gramian $P$ remains the same as in BT, while the weighted observability Gramian $Q_{\text{S}}$ solves the following Riccati equation:
@@
2.5 Quadrature-based Balanced Truncation (QuadBT) [1]
@@
$P$ $\approx\tilde{P}=\sum_{i=1}^{v}w_{p,i}^{2}(j\sigma_{i}I-A)^{-1}BB^{T}(-j\sigma_{i}I-A^{T})^{-1},$
$Q$ $\approx\tilde{Q}=\sum_{i=1}^{w}w_{q,i}^{2}(-j\mu_{i}I-A^{T})^{-1}C^{T}C(j\mu_{i}I-A)^{-1},$
where $\sigma_{i}$ and $\mu_{i}$ are the quadrature nodes, and $w_{p,i}^{2}$ and $w_{q,i}^{2}$ are their corresponding weights. The factorizations $\tilde{P}=\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $\tilde{Q}=\tilde{L}_{q}\tilde{L}_{q}^{*}$ can be expressed as: $\tilde{L}_{p}=\hat{V}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}\hat{L}_{q}$, with $\hat{V}$ and $\hat{W}$ as in (1) and (2), respectively, and
$\hat{L}_{p}$ $=\text{diag}(w_{p,1},\dots,w_{p,v})\otimes I_{m},$
$\hat{L}_{q}$ $=\text{diag}(w_{q,1},\dots,w_{q,w})\otimes I_{p}.$
@@
Finally, the ROM in QuadBT is constructed as:
@@
    Practical methods to acquire these samples are not yet available and are left for future work.
@@
2.6 Data-driven ADI-based Balanced Truncation (DD-ADI-BT) [2]
@@ 3 Main Work
In this section, we present several new results on data-driven MOR and controller design, outlined as follows:
@@ 3.1 RADI-based Non-intrusive Implementation of LQG-BT
    Given $\hat{A}=\hat{P}_{\text{LQG}}(-S_{v}^{*}+\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C})\hat{P}_{\text{LQG}}^{-1}$, the eigenvalues of $\hat{A}$ are equal to those of $S_{v}^{*}+\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$. As $S_{v}^{*}$ is Hurwitz, the matrix $\hat{A}-\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$ is Hurwitz with eigenvalues at $(-\sigma_{1}^{*},\cdots,-\sigma_{v}^{*})$.
It is clear from Theorem 3.1 that the RADI-based approximation of $P_{\text{LQG}}$ is given by $P_{\text{LQG}}\approx\hat{V}Q_{v}^{-1}\hat{V}^{*}$.
@@
Let $\hat{W}$ be as defined in (2), with all interpolation points $(\mu_{1},\dots,\mu_{w})$ located in the right half of the $s$-plane. Assume further that the pair $(S_{w},L_{w})$ is controllable and that $P_{w}>0$ uniquely solves the Lyapunov equation:
@@
Again, it is clear from Theorem 3.2 that the RADI-based approximation of $Q_{\text{LQG}}$ is given by $Q_{\text{LQG}}\approx\hat{W}P_{w}^{-1}\hat{W}^{*}$.
Similar to DD-ADI-BT [2], the RADI-based non-intrusive implementation of LQG-BT can be directly obtained using the results of Theorems 3.1 and 3.2. The pseudo-code for this implementation, named “Data-driven LQG-BT (DD-LQG-BT)”, is provided in Algorithm 1.
Algorithm 1 $D$-LQG-BT
@@
[…]